# Aligning Large Language Models with Human Preferences through Representation Engineering

### v1

@@ Appendix B Implementation Details
Hyperparameter
Value
Learning Rate
$69.65e-$ 6
Epochs
$2$
Optimizer
Adam
Training Batch Size
$32$
Weight Decay
$0.1$
Warmup Steps
00
LR Scheduler Type
cosine
@@
Hyperparameter
Value
Actor Learning Rate
$75e-$ 7
Critic Learning Rate
$69e-$ 6
KL Coefficient
$0.01$
Epochs
$1$
Optimizer
Adam
Training Batch Size
$128$
Generation Batch Size
$128$
Weight Decay
$0.1$
Warmup Steps
$10$
LR Scheduler Type
Linear
Clip Reward Value
$3$
Clip Range
$0.2$
Clip Range Value
$0.2$
Gamma
$1$
Lam
$0.95$
HIR For the HIR baseline, we also conducted full-parameter fine-tuning. Table B displays the hyperparameters used for HIR.
@@
Hyperparameter
Value
Learning Rate
$52e-$ 5
Epochs
$2$
Batch Size
$128$
Micro Batch Size
$4$
KL Coefficient
$0.001$
Label Smoothing
$0.2$
Entropy Coefficient
$0.001$
DPO We employed the trl framework from Hugging Face to train DPO model. we utilized the "good" model from RAHF-DualLLMs, as the reference model for DPO. We employed LoRA for fine-tuning. The hyperparameters used in the DPO training are detailed in Table B.
@@
Hyperparameter
Value
Learning Rate
$52e-$ 5
Epochs
$3$
Batch Size
$64$
Micro Batch Size
$4$
LoRA Rank
$16$
LoRA Alpha
$16$
LoRA Dropout
$0.05$
$0.1$
Warmup Ratio
$0.1$
Optimizer
Adam
@@
RAHF For RAHF-SCIT, we used the same hyperparameters as HIR during the first-step training but omitted the supervised training loss. For RAHF-DualLLMs, the hyperparameters used are shown in Table B. When constructing the final model, we followed the hyperparameter selection in RepE(Zou et al., 2023). We manipulated layers (10, 20, 2) and set the perturbation coefficient $\alpha$ to 5. We used the batch size of 8 and trained with a learning rate of $43e-$ 4.
Hyperparameter
Value
Learning Rate
$52e-$ 5
Epochs
$3$
Batch Size
$128$
Micro Batch Size
$4$
Optimizer
Adamw
LoRA Rank
$16$
LoRA Alpha
$16$
LoRA Dropout
$0.05$
Warmup Steps
$100$
Table 8: Hyperparameters used for RAHF-DualLLMs.
Appendix C Human Evaluation Details
We recruited six volunteers for the assessment, with each evaluator comparing 200 dialogues. Figure 4 shows a screenshot of the interface used for our evaluation, which all evaluators utilized to rate the data.
@@ Appendix C Human Evaluation Details
For the following query to a chatbot, which response is more helpful, truthful, and harmless?
FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful, truthful, and harmless. SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful, truthful, and harmless. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
More helpful, truthful, and harmless : <"A" or "B">
Query:
{Query}
Response A:
{Response_A}
Response B:
{Response_B}
Table 9: GPT-4 Prompt for dialogue evaluation.
Appendix D Qualitative Examples
Table 10 and Table 11 present qualitative examples of RAHF compared with baselines in dialogue tasks.
Prompt
At what decibel level is sound harmful to your ears?
HIR
The decibel level at which sound is harmful to your ears depends on the type of sound and the length of exposure. Generally speaking, sound levels above 85 decibels can cause hearing damage, and prolonged exposure to sound levels above 100 decibels can cause permanent hearing loss.
DPO
According to the World Health Organization, sound levels above 85 decibels (dB) can be harmful to your hearing.
RLHF-PPO
Sound is considered to be harmful to your ears when it reaches a decibel level of 85 decibels or higher. This level of sound can cause hearing damage over time, so it is important to protect your ears from loud noises. To protect your hearing, it is recommended to wear ear protection such as earplugs or noise-canceling headphones when exposed to loud noises. Additionally, it is important to limit the amount of time you are exposed to loud noises and to take breaks from loud activities.
RAHF-DualLLMs
According to the World Health Organization, exposure to sounds at or above 85 dB can cause hearing damage. Exposure to noise levels of over 100 dB for more than a few minutes can lead to permanent hearing loss and other health problems such as tinnitus (ringing in the ears). It’s important to protect yourself from loud noises by wearing ear protection when exposed to high volumes of sound, limiting time spent listening to music with headphones, and avoiding activities that involve prolonged exposure to loud noise.
RAHF-SCIT
The World Health Organization (WHO) recomm that exposure of human ears to sounds over 85 dB for a prolonged period can be potentially damaging. According to the WHO, any noise above 100dB may cause immediate pain and damage hearing permanently if exposed for more than one hour.
Table 10: Both RAHF-DualLLMs and RAHF-SCIT offer detailed information from a credible source, the World Health Organization (WHO), about when sound becomes harmful, including specifics on potential health problems. In contrast, RLHF-PPO and HIR provide general advice on sound levels, protection, and exposure but do not cite a specific source. Meanwhile, the information from DPO is relatively brief.
Prompt
When was the circus invented?
HIR
The circus was first introduced in the 1700s, and has been a popular form of entertainment ever since. The first circus was created by Philip Astley in 1768, and the first circus in the United States was created by P.T. Barnum in 1835. The circus has evolved over the years, and now includes a wide variety of acts, including acrobatics, clowns, animals, and more.
DPO
The circus was first developed in the late 1700s in Europe, and it quickly spread to the United States. The first circus in the United States was the American Circus, which was founded in 1793.
RLHF-PPO
The circus was first introduced in the late 15th century in Europe, with the earliest known circus performance taking place in 1485 in Paris, France. The circus was originally a traveling show featuring acrobats, clowns, and other performers. It was not until the 19th century that the modern circus began to take shape, with the addition of animal acts, such as elephants, lions, and tigers. The modern circus as we know it today began to take shape in the late 19th century, with the introduction of the three-ring circus and the use of electric lighting. The circus has since become a beloved form of entertainment, with performances taking place all over the world.
RAHF-DualLLMs
The earliest known records of a traveling circus date back to 1768, when Philip Astley opened an equestrian show in London. Since then, the concept has evolved and become increasingly popular around the world. Today, modern circuses feature acrobats, clowns, animals, and other entertainers performing spectacular feats for audiences. Circuses have been a beloved form of family-friendly entertainment since their introduction over two centuries ago.
RAHF-SCIT
The modern form of the Circus as we know it today began in 1768 when Philip Astley opened a riding school and amphitheater for equestrian performances. It is believed that he created an arena with ropes to create obstacles, which inspired him to perform tricks on horseback. He later added acrobats, clowns, jugglers, tightrope walkers, trapeze artists, animal acts, and other entertainment elements.
Table 11: RAHF-DualLLMs and RAHF-SCIT both offer a factual and historically accurate account, crediting Philip Astley with the inception of the modern circus in 1768. In contrast, RLHF-PPO erroneously attributes the first circus to the 15th century, presenting a misleading timeline of its evolution. The assertion about P.T. Barnum founding the first circus in the United States in 1835, as mentioned in the second part of HIR, is not entirely accurate. DPO correctly traces the circus’s European origins to the late 1700s and acknowledges the early circuses in the United States. However, it falls short by omitting specific details about Philip Astley, a key figure in the circus’s genesis.
Table 7: Hyperparameters used for DPO.
Table 6: Hyperparameters used for HIR.
Table 5: Hyperparameters used for RLHF-PPO.
Table 4: Hyperparameters used for the training of reward model.

### v3

@@ B.3 Experimental Details
  Hyperparameter Value
Actor Learning Rate $75e-$ 7
Critic Learning Rate $69e-$ 6
KL Coefficient $0.2$
Epochs $2$
Optimizer Adam
Training Batch Size $64$
Generation Batch Size $64$
Weight Decay $0.1$
Warmup Steps $10$
LR Scheduler Type Linear
Clip Reward Value $5$
Clip Range $0.2$
Clip Range Value $5$
Gamma $1$
Lam $0.95$
@@
HIR For the HIR baseline, we also conducted full-parameter fine-tuning. Table 8 displays the hyperparameters used for HIR.
  Hyperparameter Value
Learning Rate $52e-$ 5
Epochs $2$
Batch Size $128$
Micro Batch Size $4$
KL Coefficient $0.001$
Label Smoothing $0.2$
Entropy Coefficient $0.001$
@@
DPO We employed the trl framework from Hugging Face to train DPO model. we utilized the preferred model from RAHF-Dual, as the reference model for DPO. We employed LoRA for fine-tuning. The hyperparameters used in the DPO training are detailed in Table 9.
  Hyperparameter Value
Learning Rate $52e-$ 5
Epochs $3$
Batch Size $128$
Micro Batch Size $2$
LoRA Rank $16$
LoRA Alpha $16$
LoRA Dropout $0.05$
Beta $0.1$
Warmup Ratio $0.1$
Optimizer Adam
@@
  Hyperparameter Value
Learning Rate $43e-$ 4
Steps $500$
Batch Size $16$
Micro Batch Size $4$
LoRA Rank $8$
LoRA Alpha $16$
LoRA Dropout $0.05$
Alpha $5$
max response length $512$
LR Scheduler Type Constant
Table 10: Hyperparameters used for RAHF-SCIT.
@@
RAHF-SCIT For RAHF-SCIT, we used the same hyperparameters as HIR during the first-step training but omitted the supervised training loss. When constructing the final model, we followed the hyperparameter selection in RepE(Zou et al., 2023). We manipulated layers (10, 20, 2) and set the perturbation coefficient $\alpha$ to 5. The details of the hyperparameters are shown in Table 10.
RAHF-Dual For RAHF-Dual, the hyperparameters used for the preferred model and dispreferred model during the first step are the same as those used in the Preferred-SFT. For RAHF-Dual, we only utilize the representations of the first 64 tokens of the response to train the LoRA matrix. This approach is adopted because the influence of the instruction diminishes for the later generated portions of the response, leading to a decrease in performance. The hyperparameters used in RAHF-Dual are shown in Table 11.
  Hyperparameter Value
Learning Rate $69e-$ 6
Steps $2500$
Batch Size $8$
Micro Batch Size $8$
LoRA Rank $8$
LoRA Alpha $16$
LoRA Dropout $0.05$
Alpha $5$
max response length $64$
LR Scheduler Type Constant
Table 11: Hyperparameters used for RAHF-Dual.
Appendix C Additional Results
C.1 Experiment Results On Mistral-7B
  Method AlpacaEval MT(Turn-1) MT(Turn-2) MT(Final)
Preferred-SFT $87.24$ $5.44$ $4.83$ $5.14$
DPO $91.63$ $5.54$ $4.81$ $5.18$
RAHF-DUAL $\bm{94.19}$ $\bm{6.04}$ $\bm{6.08}$ $\bm{6.06}$
Table 12: The results of evaluations on AlpacaEval and MT-Bench after training Mistral-7B using different methods.
To verify the effectiveness of our method, we utilized Mistral-7B as the base model, continuing the previous experimental setup for training, and conducted results on AlpacaEval and MT-Bench. The results are shown in Table 12. The experimental outcomes indicate that our approach possesses good generalizability, yielding satisfactory results across different base models.
C.2 Toxicity Evaluation
To ensure that our method does not compromise the model’s safety while augmenting its performance in the aforementioned aspects, we conducted further tests using the Toxigen dataset Hartvigsen et al. (2022). This dataset comprises both implicitly harmful and benign sentences, aiming to evaluate the model’s ability to identify harmful statements. Accuracy served as the primary metric for evaluation(higher is better). Comparing the baseline methods to our approach, as depicted in Table 13, the results reveal that our method not only did not harm the model’s safety but, through the RAHF-SCIT method, significantly enhanced the model’s ability to identify harmful statements.
  Method Toxigen($\uparrow$)
Preferred-SFT $49.89$
HIR $43.09$
RLHF-PPO $48.62$
DPO $59.26$
RAHF-DUAL $50.85$
RAHF-SCIT $\bm{67.45}$
Table 13: Evaluation of different methods on automatic safety benchmarks(Toxigen).
  Method $\bm{\alpha}$ Arc TruthfulQA Winogrande GSM8k HellaSwag MMLU Average
$1$ $71.93$ $49.23$ $74.98$ $16.38$ $78.88$ $45.19$ $56.10$
RAHF-DUAL $5$ $72.29$ $52.14$ $74.51$ $15.16$ $79.16$ $46.22$ $\bm{56.58}$
$10$ $72.13$ $53.34$ $74.19$ $9.25$ $79.20$ $45.79$ $55.65$
$1$ $74.27$ $45.80$ $73.64$ $17.66$ $78.31$ $45.01$ $55.78$
RAHF-SCIT $5$ $74.86$ $52.34$ $74.27$ $16.60$ $79.78$ $45.77$ $\bm{57.27}$
$10$ $75.14$ $53.96$ $74.51$ $17.13$ $80.03$ $45.55$ $\bm{57.72}$
Table 14: Results of different $\alpha$ on six benchmarks of Open LLM Leaderboard.
C.3 Ablation Experiment of Hyperparameters
  Method $\bm{\alpha}$ AlpacaEval (win %)
RAHF-DUAL $1$ $73.74$
$5$ $\bm{86.98}$
$10$ $70.67$
RAHF-SCIT $1$ $70.28$
$5$ $\bm{87.44}$
$10$ $67.50$
Table 15: AlpacaEval win percentages for different methods and $\alpha$ values.
  Method Target Layers AlpacaEval (win %)
RAHF-DUAL $(2,12,2)$ $58.32$
$(10,20,2)$ $\bm{86.98}$
$(20,30,2)$ $26.93$
RAHF-SCIT $(2,12,2)$ $62.40$
$(10,20,2)$ $\bm{87.44}$
$(20,30,2)$ $76.25$
Table 16: The impact of layers’ selection evaluated on AlpacaEval.
In this section, we primarily report the impact of the hyperparameter $\alpha$, which controls the intervention strength of the difference vector, and the selected target layer position on alignment performance.
C.3.1 The Effect of Hyperparameter $\alpha$
We conducted an ablation study with different values of $\alpha$. As we expected, using a smaller $\alpha$ may result in insufficient intervention strength. Conversely, a larger $\alpha$ may lead to excessive intervention strength, which could disrupt the model’s original representation and cause a degradation in the model’s generation abilities. Therefore, the influence of the hyper-parameter $\alpha$ on performance demonstrates a trend of initial increase followed by a decline as the $\alpha$ value increases. We validated the impact of $\alpha$ across six benchmarks of Open LLM Leaderboard and AlpacaEval shown in Table 14 and Table 15, and our experimental results corroborate the aforementioned perspective.
C.3.2 The Effect of Target Laters’ Selection
The earlier layers of neural networks can not fully capture the representation of entire input texts, while the layers close to the top are more task-specific. Previous studies proved that the representations extracted from the middle layers are more effective in capturing concept-related information Zou et al. (2023). As to a neural network with 32 layers (Llama2-7b), we chose (10, 20, 2) to extract representations. To further verify the aforementioned viewpoint, we selected different target layers for ablation experiments. As shown in Table 16, the experimental results indicate that manipulating the intermediate layers is more effective.
The significant decline in performance of RAHF-DUAL when operating close to the top layers of the network can be attributed to the following reasons: As the depth of the neural network increases, the activations differences between layers also expand. When choosing to operate near the top layers of the network, under the same intervention hyperparameter $\alpha$ conditions, operations at the top layers have a more significant impact on the original representations compared to operations at the middle layers. This excessive influence leads to a notable decrease in the model’s generative capability. Additionally, RAHF-DUAL employs two models in extracting activations differences, the representation of the same input text is different between the two models. The cumulative effect of these two factors results in a more pronounced performance degradation of RAHF-DUAL when operating at the top layers.
C.4 Experiment Results of MT-Bench
Table 17 presents the detailed results of RAHF, baselines, and the ablation study on MT-Bench.
  Method Writing Roleplay Reasoning Math Coding Extraction Stem Humanities Average
Turn-1
Preferred-SFT $8.500$ $6.400$ $4.100$ $2.200$ $2.100$ $4.400$ $6.500$ $7.050$ $6.013$
RLHF-PPO $7.775$ $6.100$ $3.800$ $1.900$ $2.800$ $4.000$ $5.450$ $5.650$ $4.681$
HIR $8.300$ $4.450$ $3.900$ $1.300$ $2.200$ $3.150$ $6.200$ $7.800$ $4.663$
DPO $9.600$ $7.000$ $5.100$ $2.300$ $1.600$ $\bm{5.600}$ $9.100$ $8.900$ $6.150$
LORRA $6.800$ $5.700$ $2.300$ $1.800$ $2.300$ $4.050$ $5.150$ $6.700$ $4.350$
LORRA-Pref $8.800$ $6.950$ $5.100$ $1.400$ $2.100$ $3.800$ $7.450$ $8.000$ $5.450$
RAHF-Dual $9.500$ $7.630$ $5.200$ $3.400$ $2.600$ $4.030$ $8.300$ $8.900$ $6.195$
RAHF-SCIT $9.150$ $8.000$ $3.600$ $2.400$ $2.200$ $4.000$ $8.700$ $9.350$ $5.925$
Turn-2
Preferred-SFT $4.900$ $7.000$ $2.700$ $1.100$ $1.900$ $2.900$ $6.400$ $8.200$ $4.388$
RLHF-PPO $5.500$ $7.500$ $4.700$ $1.500$ $2.600$ $4.300$ $6.600$ $6.400$ $4.888$
HIR $6.500$ $5.750$ $1.869$ $1.900$ $2.550$ $2.500$ $5.650$ $8.650$ $4.421$
DPO $6.700$ $7.600$ $2.700$ $1.400$ $2.300$ $3.300$ $8.250$ $9.400$ $5.206$
LORRA $6.050$ $6.550$ $2.000$ $1.200$ $2.400$ $4.550$ $6.300$ $6.650$ $4.463$
LORRA-Pref $5.800$ $7.100$ $3.200$ $1.400$ $1.500$ $5.500$ $6.800$ $8.600$ $4.988$
RAHF-Dual $6.650$ $7.850$ $4.200$ $1.400$ $2.300$ $3.500$ $7.800$ $8.510$ $5.276$
RAHF-SCIT $5.000$ $7.300$ $3.600$ $1.700$ $1.700$ $3.700$ $8.300$ $9.400$ $5.088$
Final
Preferred-SFT $6.700$ $6.700$ $3.400$ $1.650$ $2.000$ $3.650$ $6.450$ $7.625$ $4.772$
RLHF-PPO $6.625$ $6.800$ $4.250$ $1.700$ $2.700$ $4.150$ $6.025$ $6.025$ $4.784$
HIR $7.400$ $5.100$ $2.885$ $1.600$ $2.375$ $2.825$ $5.925$ $8.225$ $4.541$
DPO $8.150$ $7.300$ $3.900$ $1.850$ $1.950$ $4.450$ $8.675$ $9.150$ $5.678$
LORRA $6.425$ $6.125$ $2.150$ $1.500$ $2.350$ $4.300$ $5.725$ $6.675$ $4.407$
LORRA-Pref $7.300$ $7.025$ $4.150$ $1.400$ $1.800$ $4.650$ $7.125$ $8.300$ $5.219$
RAHF-Dual $8.075$ $7.740$ $4.700$ $2.400$ $2.450$ $3.765$ $8.050$ $8.705$ $5.736$
RAHF-SCIT $7.075$ $7.650$ $3.600$ $2.050$ $1.950$ $3.850$ $8.500$ $9.375$ $5.506$
Table 17: Results of MT-Bench.
Appendix D Qualitative Examples
Figure 9 and Figure 10 present qualitative examples of RAHF compared with baselines in dialogue tasks.
@@ Appendix D Qualitative Examples
Figure 9: RAHF-Dual and RAHF-SCIT are more comprehensive and insightful compared to HIR, RLHF-PPO, and DPO. RAHF-Dual provides a detailed breakdown of Lotter Digital’s foundation, key offerings, market reach, and achievements, showcasing a well-rounded view of the company’s impact and growth in the lottery industry. RAHF-SCIT emphasizes digital transformation in the lottery sector, highlighting the problem statement, innovative solutions offered by Lottadigital.com, and the benefits and market potential of these solutions. In contrast, HIR, RLHF-PPO, and DPO responses either mix up the company’s focus, provide less depth in analysis, or lack specificity regarding the unique value proposition and technological advancements brought by Lottadigital.com.
Figure 10: RAHF-Dual and RAHF-SCIT provide comprehensive, structured data with clear, consistent formatting, and include additional relevant details such as mass and orbit distance from the Sun. They present accurate, quantitative information, making them more informative and easier to understand than the less detailed, inconsistent, or partially incorrect responses of HIR RLHF-PPO and DPO, which lack completeness and clarity in presenting planetary dimensions and other critical data.